if sin(theta)=7/25 and theta in second quadrant find the value of sec( theta)+tan(theta)?
Answer by lwsshak3(11628)

if sin(theta)=7/25 and theta in second quadrant find the value of sec(theta)+tan(theta)?
***
sinx=7/25
cosx=√(1-sin^2(x))=√(1-49/625)=√(576/625)=-24/25
secx=1/cosx=-25/24
tanx=sinx/cosx=-7/24
secx+tanx=-25/24-7/24=-32/24=-4/3